A airplane is flying at 300m/s(627m/h).How much time must elapse before a clock in the airplane and one on the ground differ by 1.00s  

Grade:12th pass

1 Answers

Umar Farooq
5 years ago
A moving clock ticks slower by \frac{t{}`}{t}= \sqrt{1-\frac{v^2}{c^2}} = 0.9999999999999999
where t is stationary time. For 1 second difference, t-t{}` = 1
t- 0.9999999999999999t=1
t=10^{^{16}}seconds=317097919.8376458883 years
t(1- 0.9999999999999999)=1
t=1/